# Force Fields for Deep Eutectic Mixtures: Application to Structure and 2D-Infrared Spectroscopy

### v1

@@ Abstract
si University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Zurich]Department of Chemistry, University of Zürich, CH-8000 Zürich, Switzerland. University of Zurich]Department of Chemistry, University of Zürich, CH-8000 Zürich, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.
@@ 1 Introduction
fields”) are computationally advantageous to evaluate. Propagating
the dynamics based on Newton’s equations of motion allows to reach
long time scales for large molecular systems, e.g. sub-microsecond
for atoms.2 Such molecular dynamics (MD)
simulations have been used to investigate processes ranging from
protein folding,3 ligand binding,4
crowding in cellular environments5 to characterizing
spectroscopic properties of solutes and peptides and reactions in the
gas phase and in solution.6 However one of the
challenges remains to develop energy functions that retain the
precision of the QM methods they are often based on.
@@
systems.7, 8 The molecular-level understanding
@@
approach.9, 10 The atomistic simulations
@@
terms.11 For applications to spectroscopy the
chemical bonds (stretching vibration) need to be described somewhat
more realistically to account for mechanical anharmonicity for which
Morse-oscillators are often sufficient because experiments at ambient
temperatures are not sensitive to highly vibrationally excited
states. Nevertheless, there is scope to use more accurate
representations, for example based on machine learning-type
approaches, in particular if reference data from high-level electronic
structure calculations are available. For the nonbonded interactions -
which include electrostatic and van der Waals terms - more
physics-based models that go beyond the standard representations have
been developed.
@@
density.12 Including higher-order atomic multipoles
@@
further improvements of the computational models.17 From
@@
parametrization18 and modified combination
rules19, 20
Deep eutectic mixtures (DES) are electrostatically demanding systems.
@@ 2.1 Simulation Setup
program.21 The molar composition of the systems
@@
9 system compositions were set up using PACKMOL.24
@@
thermostat.25 For each system composition a total of
ns were sampled.
@@ 2.2 Inter- and Intramolecular Interactions
(CGenFF)26, the corresponding TIP3P water
model27 to be used together with CGenFF, and
@@
work28, 29 and acetamide was described by
@@
(RKHS)30, 31 representation based on ab
@@
using the MOLPRO program package.32, 33 The
@@
and for different angles are shown in Figure S1.
@@
(fMDCM).34 For this, 8 point charges were
@@
using the Gaussian suite of programs.35 The grid
@@
Refs. 36, 37. Starting from this model as
@@
minimization method38 was used to optimize
@@
literature,39 the diagonal elements of computed
@@
respectively.40, 41 Thus, a reduced
@@
(MTP)14, 42, 43 representation of
SCN- used in previous work was used as well.10
@@
parameters.44 Three models were considered in the
@@
parameters originally from Bian et al.28 and
@@
parameters from Tesei et al.,29 and model M2
@@
in the literature (Ref. 28 for M0 and M2
and Ref. 29 for M1) and for model M2 the
LJ parameters from Bian and coworker were chosen for
K+.28
@@ 2.3 Analysis
Hydration free energies for the SCN- anion in water solvent were computed from thermodynamic integration.9 One SCN- anion was sampled in the gas phase and in pure water. The condensed-phase simulations were carried out in the NpT ensemble with 997 water molecules (cubic box size Å3).45, 9 Simulations with 24 different coupling parameters for electrostatic and vdW interactions, respectively, were performed for each a gas phase and solvated SCN-. Initial conditions for these simulation were taken from an unbiased simulation, equilibrated for 50 ps with the respective coupling parameter and run for another 150 ps for statistical sampling. In general, the hydration free energy is accumulated from
@@
nonbonded interaction exclusion.46 Therefore, only
@@
For computing the frequency fluctuation
correlation function (FFCF), the frequency trajectories
for each oscillator are required. These were determined from an
instantaneous normal mode (INM)47 analysis for the CN
vibrational frequencies for all SCN- ions
on snapshots of the trajectories every fs along
the first 2 ns of each production simulation (10 ns in total for
each system composition).
For the INM analysis the structure of every SCN- ion was optimized
while freezing the position of all remaining atoms in the system,
followed by a normal mode analysis using the same force field that was
employed for the MD simulations. Previously, such an approach has been
validated for N in solution by comparing with rigorous quantum
bound state calculations.48
From the frequency trajectory for each oscillator the FFCF was determined which contains information on relaxation time scales corresponding to the solvent dynamics around the solute. The FFCFs were fit to an empirical expression
using an automated curve fitting tool from the SciPy
library.49 Here, , , and
@@
The experimental data shown in this paper (Figure 7) originate from the same raw data as those of Ref. 10, this time extracting the FFCF from the 2D-IR spectra. In brief, a 2D IR spectrometer in pump-probe with scatter suppression by quasi-phase-cycling has been used to obtain the raw data.50 A 50%/50% mixture of two isotopologues of the salt, KS12C14N and KS13C15 has been investigated. Solutions were prepared with constant total salt concentration of 4.4 M, and the molar ratio of D2O vs acetamide varied from 0% to 100% in steps of 10%.
To extract the essential information from the 2D IR spectra, i.e., the FFCF, each 2D IR peak was modelled as a correlated 2D-Gaussian function:
where is an amplitude, the center frequencies along
frequency axis , the frequency width, and a
correlation coefficient.51 In total 8 such peaks were
defined for a 2D IR spectrum, with many parameters set equal to
minimize their total number in the global fitting. That is, two pairs
@@
solvent environment. Each diagonal peak consists of ground state
bleach (GB)/stimulated emission (SE) as well as excited state
absorption (EA), where the latter has been down-shifted in the
@@
same. In addition, two cross peak pairs were introduced, each again
consisting of GB/SE and EA contributions, for which center frequencies
and width were assumed to be the same as for the corresponding
diagonal peaks along the two frequency axes. The amplitudes of the
two cross peak pairs were set to be the same, and was assumed
for the correlation, since cross-relaxation randomizes the
frequency.
The property of interest here is the frequency correlation of the diagonal peaks as a function of the waiting time between the pump-pulse pair and the probe pulse. For times larger than a few 100 fs, reflects the FFCF and can be compared to results from MD simulations after proper scaling, while dynamics faster than this is hidden in the homogeneous dephasing time .8
@@ 3.1 Adaptation and Validation of the LJ Parameters
available53 which compares with related anions such as
@@
had been used for modeling aqueous SCN- solutions.29
@@
with the experimentally measured52 density
@@
For model M2, also based on the fMDCM representation for the
SCN- anion, a different approach was taken. Here, the LJ parameters
and were adjusted to best match
counterpoise-corrected interaction energies from electronic structure
calculations. The systems considered included on SCN- anion
surrounded by (i) 16 water molecules; (ii) 14 water molecules with
either 1 SCN- or (iii) 1 K+ ion; (iv) 12 water molecules with
both 1 SCN- and K+ ion. For each type of system (i) to (iv) 50
independent conformations were extracted randomly from previous MD
simulations.10 Reference interaction energies were
then determined at the M06-2X/aug-cc-pVTZ level of theory using
Gaussian program.35 Higher levels of quantum chemical
theory, such as coupled cluster energies, are not feasible due to the
unfavourable scaling of the computations with system and basis set
size.
@@ 3. In a second step, the optimized of
experiment is found for which yields
@@ 3.2 Structure and Ordering of the Mixture
solution.56 It is important to stress that such pair
@@
distribution functions.56 Figure 4
@@
black line) are possible.56 For model M0, the
@@
increases - i.e. the ion recruits water molecules as the water density
decreases. The shift of the first maximum for model M2 relative
to model M1 is consistent with the larger values
for the S- and N-atoms for M2 compared with M1, see Tables
S3 and S4.
@@
maximum. Compared with that, models M1 and M2 correctly
capture the peak height, which is a proxy for the interaction
strength, whereas the position of the first maximum is in general
shifted to shorter separations. This, however, changes with decreasing
water content for which the S–N and N–N pair correlation functions
shift to longer separations. It should be noted that the peak position
of the anion-anion pair distribution functions in simulations can be
steered by adjusting the LJ parameter of the SCN-
atoms individually. However, the procedure followed here was not to
match pair distribution functions as the LJ parameters of
SCN- are at first either taken from literature (M0, M1)
or fit to ab initio data (M2) and later scaled
collectively to best match the density of an aqueous KSCN system or
hydration free energy of SCN- in water.
The comparison of the between atoms of different SCN- anions
in simulations using the three models for the intermolecular
interactions are shown in Figures 4D to F. Peak
intensities for SCN- atom pair in different water/acetamide solvent
mixture remains either similar (M1) or increase for larger
acetamide content (M2). For simulations with multipoles (model
M0), the peak positions are generally closer those reported from
experiments but with higher intensities which are highest for pure
water solvent and decreases for larger acetamide content in the
water/acetamide solvent mixture.
Another view on the structural SCN- radial-angular pair
distribution is shown in Figure 5 along the C-C bond
distances and (CN)-(CN) bond angles (similar to SCN- dipole
vectors). Even after adjusting the LJ parameters for M0 to yield
improved thermodynamics, the MTP-based model lead to mainly angled
SCN- pairs in pure acetamide and solvent mixtures with large
acetamide content. This changes to predominantly parallel and, partly,
@@
unmodified MTP-based model.10 In simulations using
fMDCM for the electrostatics (models M1 and M2), the
@@
solvent mixtures and are close to squared cosine angle distributions
which are statistically expected for weakly or non-interacting
particles. This can be even better seen in Figure 6
which reports the 2-dimensional projections of the distributions shown
in Figure 5 using model M2.
@@
that appear for acetamide as the solvent disappear as the water
content increases.
@@ 3.3 Dynamics of the Mixtures
mixtures. Assuming a single exponential significantly deteriorates the
fit. To allow direct comparison between experiment and simulations,
the experimental data was fit to a double-exponential decay because
for the computed FFCFs a second, short-time decay (to which the
experiments are insensitive) needs to be included for fitting the full
time range of the data, see Figures S2A to C. The
set of optimized parameters of the bi-exponential fit function are
shown in Figure S3 and Table S6.
Qualitatively, models M1 and M2 are rather successful in
capturing the experimentally observed FFCFs and their dependence on
the amount of water in the solvent mixture whereas model M0 is
not. Overall, the decay from simulations using M0 is too
slow. However, the relative amount by which the time constant changes
between 0% and 100 % water is considerably closer to the
experimentally observed change than in the other two models.

### v2

# Force Fields for Deep Eutectic Mixtures: Application to Structure, Thermodynamics and 2D-Infrared Spectroscopy

@@ Abstract
University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland. University of Zurich]Department of Chemistry, University of Zürich, CH-8000 Zürich, Switzerland. University of Zurich]Department of Chemistry, University of Zürich, CH-8000 Zürich, Switzerland. University of Basel]Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.
@@ 1 Introduction
fields”) are computationally advantageous to evaluate and are being
used in molecular dynamics (MD) simulations which solve the
Newtonian equations of motions. MD simulations have been used to
investigate processes ranging from protein folding,2
ligand binding,3 crowding in cellular
environments4 to characterizing spectroscopic
properties of solutes and peptides and reactions in the gas phase and
in solution.5 However one of the challenges remains
to develop energy functions that retain the precision of the QM
methods they are often based on.
@@
systems.6, 7 The molecular-level understanding
@@
approach.8, 9 The atomistic simulations
@@
terms.10 For applications to vibrational
spectroscopy, going beyond the harmonic approximation to describe
chemical bonds is necessary to include effects of mechanical
anharmonicity. Such improvements can, e.g., be achieved through the
use of machine learning-based approaches, in particular if
reference data from high-level electronic structure calculations are
available.11, 5, 12, 13, 14
For the nonbonded interactions - which include electrostatic and van
der Waals terms - more physics-based models that go beyond the
standard representations have been developed.
@@
density.15 Including higher-order atomic multipoles
@@
further improvements of the computational models.20 From
@@
parametrization21 and modified combination
rules22, 23
Deep eutectic solvents (DESs) are multicomponent mixtures
comprising hydrogen bond acceptors (HBAs) and hydrogen bond donors
(HBDs) at particular molar
ratios.24, 25, 26 In
general, DESs are characterized by a pronounced depression of their
melting point compared to their components and remain in the liquid
phase over a wider temperature range.27 If the
mixtures contain ions, the intermolecular interactions involve
pronounced electrostatic contributions which is also - in part - due
to crowding. The particular mixture considered here consists of
water, acetamide and KSCN which is present as solvated K+ and
SCN- (thiocyanate) ions.28, 29 Acetamide
forms low-temperature eutectics with a wide range of inorganic salts
and the resulting non-aqueous solvents have a high ionicity. Such
mixtures have also been recognized as excellent solvents and molten
acetamide is known to dissolve inorganic and organic compounds. The
SCN- anion is an ideal spectroscopic probe because the CN-stretch
vibration absorbs in an otherwise empty region of the infrared
region. Previous work has taken advantage of these spectral features to
probe the
effect of water addition to urea/choline chloride and in
acetamide/water mixtures.30, 9
Traditionally, energy functions for DES have been optimized by
starting from a conventional energy function and readjusting
specifically partial charges and in
some cases van der Waals parameters.31, 32, 33, 34, 35, 36, 37, 38
@@ 2.1 Simulation Setup
program.39 The molar composition of the systems
@@
9 system compositions were set up using PACKMOL.42
@@
thermostat within the extended system constant pressure and
temperature algorithm as implemented in CHARMM.43, 44
The mass of the pressure piston and piston collision frequency were
set to u and 5 ps-1, respectively, and the mass of the
thermal piston to kcal/mol ps2.
For each system composition a total of ns were sampled.
@@ 2.2 Inter- and Intramolecular Interactions
(CGenFF)45, the corresponding TIP3P water
model46 to be used together with CGenFF, and
@@
work47, 48 and acetamide was described by
@@
(RKHS)49, 50 representation based on ab
@@
using the MOLPRO program package.51, 52 The
@@
and for different angles are shown in Figure
S1.
@@
(fMDCM).53 For this, 8 point charges were
@@
using the Gaussian suite of programs.54 The grid
@@
Refs. 55, 56. Starting from this model as
@@
minimization method57 was used to optimize
@@
literature,58 the diagonal elements of computed
@@
respectively.59, 60 Thus, a reduced
@@
(MTP)17, 61, 62 representation of
SCN- used in previous work was used as well.9
@@
parameters.63 Three models were considered in the
@@
parameters originally from Bian et al.47 and
@@
parameters from Tesei et al.,48 and model M2
@@
in the literature (Ref. 47 for M0 and M2
and Ref. 48 for M1).
For model M2, also based on the fMDCM representation for
the SCN- anion, a different approach was taken. Here, the LJ
parameters and were adjusted to best
match counterpoise-corrected64 interaction energies
from electronic structure calculations. The systems considered
included one SCN- anion surrounded by (i) 16 water molecules;
(ii) 14 water molecules with one additional SCN-, (iii)
14 water molecules with one K+ ion; (iv) 12 water molecules
with both one additional SCN- and K+ ion. For each type
of system (i) to (iv) 50 independent conformations were extracted
randomly from previous MD simulations.9
Reference interaction energies were then determined at the
M06-2X/aug-cc-pVTZ level of theory using Gaussian
program.54 Higher levels of quantum chemical theory,
such as coupled cluster energies, are not feasible due to the
unfavourable scaling of the computations with system and basis set
size.
@@ 2.3 Analysis
Hydration free energies for the SCN- anion in water solvent were computed from thermodynamic integration.8 One SCN- anion was sampled in the gas phase and in pure water. The condensed-phase simulations were carried out in the NpT ensemble with 997 water molecules (cubic box size Å3).65, 8 Simulations with 24 different coupling parameters for electrostatic and vdW interactions, respectively, were performed for each a gas phase and solvated SCN-. Initial conditions for these simulation were taken from an unbiased simulation, equilibrated for 50 ps with the respective coupling parameter and run for another 150 ps for statistical sampling. In general, the hydration free energy is accumulated from
@@
nonbonded interaction exclusion.66 Therefore, only
@@
For computing the frequency fluctuation correlation function (FFCF),
the frequency trajectories for each oscillator are
required. These were determined from an instantaneous normal mode
(INM)67 analysis for the CN vibrational frequencies
for all SCN- ions on snapshots of the trajectories every
fs along the first 2 ns of each production simulation (10 ns
in total for each system composition). For the INM analysis the
structure of every SCN- ion was optimized while freezing the
position of all remaining atoms in the system, followed by a normal
mode analysis using the same force field that was employed for the MD
simulations. Previously, such an approach has been validated for
N in solution by comparing with rigorous quantum bound state
calculations.11
From the frequency trajectory for each oscillator the FFCF was determined which contains information on relaxation time scales corresponding to the solvent dynamics around the solute. The FFCFs were fit to an empirical expression68, 8
using an
automated curve fitting function (scipy.optimize.curve_fit)
from the SciPy library using the default trust region reflective
algorithm.69
Here, , , and
@@
The experimental data shown in this paper (Figure 7)
originate from the same raw data as those of
Ref. 9, this time extracting the FFCF from the
2D-IR spectra. Technical details of the 2D IR spectrometer used
to obtain the raw data are described in Ref. 70. A
50%/50% mixture of two isotopologues of the salt, KS12C14N
and KS13C15N has been investigated. Solutions were
prepared with constant total salt concentration of 4.4 M, and the
molar ratio of D2O vs acetamide varied from 0% to 100% in steps
of 10%.
Very generally speaking, the 2D IR spectrum of two coupled modes consist of two diagonal contributions for each mode, and two cross peak contributions that appear when energy transfer occurs between both modes. Furthermore, each one of these contributions consist of a pair of a negative and positive peak, i.e., ground state bleach and excited state absorption, respectively. Hence, eight peaks exist in total.7 To extract the essential information from the 2D IR spectra, i.e., the FFCF, each one of these 2D IR peaks was modelled as a correlated 2D-Gaussian function:
where is an amplitude, the center frequencies along frequency axis , the frequency width, and a correlation coefficient.71 In the limiting case of , the 2D Gaussian is uncorrelated and a circular peak would appear. For the 2D Gaussian is fully correlated with the peak stretched along the diagonal and widths in the diagonal direction and 0 in the anti-diagonal direction. To minimize the total number of parameters in the global fitting, many of the parameters for modeling the eight peaks as Gaussian functions were constrained to be equal.
That is, two pairs
@@
solvent environment. For each diagonal contribution, the ground state
bleach was positioned on the diagonal (i.e., ), while the excited state absorption has been down-shifted in the
@@
same. The center frequencies and widths of the cross peak pairs were assumed to be the same as for the corresponding diagonal peaks. The
amplitudes of the two cross peak pairs were set to be the same,
and was assumed for the correlation, since cross-relaxation
randomizes the frequency.
The property of interest here is the frequency correlation of the diagonal peaks as a function of the waiting time between the pump-pulse pair and the probe pulse. It evolves from ideally 1 at to 0 for , however, dynamics faster than few 100 fs is hidden in the homogeneous dephasing time and cannot be observed.7 Beyond a few 100 fs, reflects the normalized FFCF and can be compared to results from MD simulations after proper scaling.
@@ 3.1 Adaptation and Validation of the LJ Parameters
available73 which compares with related anions such as
@@
had been used for modeling aqueous SCN- solutions.48
@@
with the experimentally measured72 density
@@ 3. In a second step, the optimized of
experiment was found for , very close to
(no modifications in the van der Waals parameters), which yields
@@ 3.2 Structure and Ordering of the Mixture
solution.76 It is important to stress that such pair
@@
distribution functions.76 Figure 4
@@
black line) are possible.76 For model M0, the
@@
increases. In other words, with decreasing water density the ion
recruits water molecules. The shift of the first maximum for model
M2 relative to model M1 is consistent with the larger
values for the S- and N-atoms for M2 compared with
M1, see Tables S3 and
S4.
@@
maximum. Compared with that, models M1 and M2
correctly capture the peak height for and , which is a proxy for the interaction strength. The
position of the first maximum, however, is at larger separation as
seen for model M0. Opposite to the reference EPSR results,
the position of the first maximum in is at larger
separation than in due to a higher concentration of
negative charge on the N- rather than the S-atom in SCN-
according to the fitted fMDCM ( e
vs. e) and MTPL ( e vs. e) models.
It should
be noted that the peak position of the anion-anion pair distribution
functions in simulations can be steered by adjusting the LJ parameter
of the SCN- atoms individually. However, the
procedure followed here was not to match pair distribution functions
as the LJ parameters of SCN- are at first either taken from
literature (M0, M1) or fit to ab initio data (M2) and later scaled collectively to best match the density of an
aqueous KSCN system or hydration free energy of SCN- in water.
Peak intensities for SCN- atom pairs in different water/acetamide
solvent mixture remains either similar (M1) or increase for
larger acetamide content (M2). For simulations with multipoles
(model M0), the peak positions are generally closer those
reported from experiments but with higher intensities which are
highest for pure water solvent and decreases for larger acetamide
content in the water/acetamide solvent mixture.
Figure 5 reports radial-angular pair distribution
functions for SCN-/SCN- pairs along the C-C bond distances
and (CN)-(CN) bond angles (related to SCN- dipole vectors). Even
after adjusting the LJ parameters for M0 to yield improved
thermodynamics, the MTP-based model lead to mainly angled SCN-
pairs in pure acetamide and solvent mixtures with large acetamide
content. This changes to predominantly parallel and, partly,
@@
unmodified MTP-based model.9 In simulations
using fMDCM for the electrostatics (models M1 and M2), the
@@
solvent mixtures and are close to a sine-angle distribution
which is the probability to find a particle at a certain
angle on a sphere ().
This can be even better seen in
Figure 6 which reports the 2-dimensional projections
of the distributions shown in Figure 5 using model M2.
@@
that are present with acetamide as the solvent disappear with
increasing water content.
@@ 3.3 Dynamics of the Mixtures
mixtures. Assuming a single exponential significantly
deteriorates the fit. To allow direct comparison between experiment
and simulations, the experimental data was also fit to a
double-exponential. The set of optimized parameters of the
bi-exponential fit function are shown in Figure S3
and Table S6. Qualitatively, models M1
and M2 capture the experimentally observed FFCFs and their
dependence on the amount of water in the solvent mixture whereas
model M0 does not. Overall, the decay on the longer
time scale from simulations using M0 is too slow. However,
the relative amount by which the time constant changes between 0% and
100 % water is considerably closer to the experimentally observed
change than in the other two models.